When twice the reciprocal of 3 times a number is subtracted from 3 the result is 7 times the reciprocal of 3 times the number. Find the number.
Can you please help me out? Thanks so much in advance:)

Answer by mananth(16946)

When twice the reciprocal of 3 times a number is subtracted from 3 the result is 7 times the reciprocal of 3 times the number. Find the number.

number -x
3 times the number = 3x
reciprocal = 1/3x
twice = 2/3x
3-(2/3x)= 7(1/3x)
(7/3x)+(2/3x)=3
9/3x=3
9x=9
x=1